OpenStudy (kaederfds):

A projectile is fired horizontally from a gun that is 45.0 m above flat ground, emerging from the gun with a speed of 250 m/s. (a) How long does the projectile remain in the air? (b) At what horizontal distance from the firing point does it strike the ground? (c) What is the magnitude of the vertical component of its velocity as it strikes the ground?

OpenStudy (kaederfds):

t=3 s and then?

OpenStudy (abhisar):

How t=3s ?

OpenStudy (kaederfds):

45=0+(1/2)(10)(t^2)

OpenStudy (abhisar):

yes that's correct !

OpenStudy (abhisar):

Now, we will have to calculate the horizontal distance covered in the same time i.e 3 seonds

OpenStudy (abhisar):

Just multiply 250 with 3 i.e speed with time.

OpenStudy (abhisar):

Got it ?

OpenStudy (kaederfds):

(b) 750 m

OpenStudy (abhisar):

yes, Now for the third part we will have to calculate the final vertical velocity. u=0 s=45 a=9.8 Calculate v by using \(\sf v^2=u^2+2as\)

OpenStudy (abhisar):

\(\sf v^2=0+2\times 9.8\times 45\\v^2=882\\v=\sqrt{882}=29.69 m/s\)

OpenStudy (kaederfds):

i got 30 m/s :) i used g=10m/s^2

OpenStudy (kaederfds):

anyway, thanks

OpenStudy (abhisar):

It's a good practice to use g=10 [if not stated]

OpenStudy (abhisar):

\(\rlap{\color{blue}{\huge\bigstar}\huge \color{blue}{ \text{You're Most Welcome! }}\color{blue}\bigstar}{\; \color{aqua}{\huge\bigstar}\huge \color{aqua}{\text{You're Most Welcome! }}\color{aqua}\bigstar}\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)